OpenStudy (anonymous):

please show me how to solve for x (x/7)=(x/8)+(10/7)

OpenStudy (anonymous):

Okay, so what we need to do here is make all our fractions have the same denominator first. Then we can just factor them out. 7 and 8 both multiply to 56, so we'll use that. 8x / 56 = 7x / 56 + 80 / 56. In each case, we multiplied the denominator so it equalled 56, then multiplied the numerator by the same amount. These are still the same fractions; if I divide 8x / 56 by 8, I get x / 7 again. So I haven't really done anything, mathematically speaking. But now we can divide both sides of our equation by 56, to remove the fractions altogether. 8x = 7x + 80. Then we just do some simple algebra, removing 7x from both sides: x = 80. And we're done!

OpenStudy (anonymous):

by simplification we get (8x/56)=(7x/56)+(80/56) multiplying both sides by 56 8x=7x+80 adding -7x on both sides 8x-7x=7x-7x+80 x=0+80 x=80 thus the value of x is 80